Is (-1,-4) on the graph of y-2x=-2?

Thank you

Answer by stanbon(57341)

Is (-1,-4) on the graph of y-2x=-2?
Substitute the x and y values to see if you get a true statement.
-4 -2(-1) = -2
-4+2 = -2
-2 = -2
It is true so the point is no the graph.
--------
Cheers,
Stan H.